OpenStudy (anonymous):

If 1/2x +1/2(1/2x + 1/2(1/2x +1/2(1/2x + ... = y, then x = ?

OpenStudy (anonymous):

\[\frac{1}{2}x +\frac{1}{2}(\color\red{\frac{1}{2}x + \frac{1}{2}(\frac{1}{2}x +\frac{1}{2}(\frac{1}{2}x + ... }= y\]red part is \(y\)